Tyler McGinnis invested some money at 3% simple interest and some at 5% simple interest. The total amount of his investment was $28,000, and the interest he earned during the first year was $1240. How much did he invest at each rate?
Answer by mananth(16946)

Part I 3.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
28000
Interest----- 1240.00

Part I 3.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 28000 -------------1
Interest on both investments
3.00% x + 5.00% y= 1240
Multiply by 100
3 x + 5 y= 124000.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -84000.00
Add this to (2)
2 y= 40000
divide by 2
y = 20000 - x)
Part I 3.00% $ 8000
Part II 5.00% $ 20000

CHECK
8000 --------- 3.00% ------- 240.00
20000 ------------- 5.00% ------- 1000.00
Total -------------------- 1240.00